OpenStudy (quickstudent):

How do I factor x^2 + 12x + 0 using the guess and check method?

OpenStudy (ujjwal):

take x common and that's it, i guess..

OpenStudy (anonymous):

x(x+12)=0 x=0 and x=-12

OpenStudy (quickstudent):

I don't understand either answer.

OpenStudy (quickstudent):

I know how to factor using guess and check method, but I don't understand how to do it to this one. I know that I have to find a factor pair of the second term that would add up to the third term, but I can' do that here. So how do I factor this using guess and check method? x^2 + 12x + 0

OpenStudy (quickstudent):

@ParthKohli ?

OpenStudy (quickstudent):

@paki ?

OpenStudy (paki):

what about quadratic formula... ?

OpenStudy (quickstudent):

What is the quadratic formula? Well anyway I have to use the guess and check method.

OpenStudy (quickstudent):

Will someone please help?

OpenStudy (larseighner):

You have had two perfectly good answers. If you can factor, as you say, just factor x out of the original equation. Lose the zero. It is is just window dressing.

OpenStudy (quickstudent):

I don't understand how to do that.

OpenStudy (quickstudent):

How do I factor x?

OpenStudy (quickstudent):

I only know how to factor constants.

OpenStudy (larseighner):

Okay \( x^2 = x \cdot x\) so your original expression is \( x \cdot x + 12x\)

OpenStudy (larseighner):

The only constant is 0. It is nothing. So just drop it.

OpenStudy (quickstudent):

Ok. I understand that, but how would I use the guess and check method for this? I'm supposed to get an answer that looks like this: (x + a)(x + b)

OpenStudy (larseighner):

Factor the x out: \[ x\cdot x + 12x = x(x+12) \] If it makes you feel better, you can rewrite x as x + 0, but it is kind of pointless. \[ x\cdot x + 12x = x(x+12) = (x+0)(x+12)\]

OpenStudy (quickstudent):

Oh, now I understand.

OpenStudy (quickstudent):

\(\color{red}{\huge\bigstar}\huge\text{Thanks!}\color{red}\bigstar\) \(~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (larseighner):

S'k